OpenStudy (anonymous):

The graph of f(x)=1/(x^2-c) has a vertical asymptote at x=3. Find c.

myininaya (myininaya):

This means the function shouldn't exist at x=3. So since you have a fraction and fractions don't exist when the bottom is zero, then when is your bottom 0?

myininaya (myininaya):

Or when is x^2-c=0 for x=3.

myininaya (myininaya):

Replace x there with 3 and then solve for c.

OpenStudy (anonymous):

so c=9?

myininaya (myininaya):

Perfect. :) x^2-9 is zero when x=3 which means the function will not exist at x=3 (or even x=-3) but our objective is completed Good job!

OpenStudy (anonymous):

Thank you!